2.00 cm = ________nm
1
Expert's answer
2020-10-02T14:19:35-0400

QUESTION # 136332

2.00 cm = ________nm

ANSWER

1cm=10-2m                 

1nm=10-9m

2cm(10-2m/1cm) × (nm/10-9m)

           = (2×10-2)/10-9nm

           =2.00×107nm